8.0 mL of 0.10 M HC2H3O2 is titrated with 0.10 M NaOH. Perform the following calculations & Show ALL work


  1. pH of the acid before any NaOH is added
  2. Volume of 0.10 M NaOH needed to reach the equivalence point
  3. pH at the equivalence point
  4. Volume of 0.10 M NaOH needed to reach the half-way point
  5. pH at the half-way point
1
Expert's answer
2021-04-27T07:49:40-0400

"PH=- log [H^+]"


"PH=-log[0.10]=1"



NaOH (aq) + HCl (aq) → NaCl (aq) + H2(l)


Volume of NaOH = 8.0 ml




pH of half-way point.


8 mL of 0.1 M acetic acid = 0.1 M / 1000 mL x 8.0 mL = 8 x 10-4 moles.

At the half-way point, 1/2 the original number of moles has been titrated.

As the molarity of the NaOH is also 0.1 M, this means 4.0 mL of NaOH has been added. Also, only 4 x 10-4 moles if acetic acid remains and 4 x 10-4 acetate has been produced.

Concentration of acetic acid = 4 x 10-4 moles / 12 mL x 1000 mL/L =

3.33 x 10-2M. As it is the half-way point, the acetate concentration is also 3.33 x 10-2 M.


According to the Henderson-Hasselbalch equation,


pH = pKa + log [acetate] /[acetic acid] = pKa + log 1 = pKa.


In this case the pH is equal to the pKa which is 4.75